# Uncovering Systemic and Environment Errors in Autonomous Systems Using Differential Testing

### v1

@@ Introduction
Autonomous agents are increasingly deployed in complex real-world applications such as autonomous driving (Yurtsever et al. 2020), crop fertilization (Gautron et al. 2022; Solow, Saisubramanian, and Fern 2025), and elderly care (Bardaro, Antonini, and Motta 2022; Mhlanga 2024). Agents operating in complex settings may sometimes produce undesirable behaviors, including failure to complete the task. We refer to such behaviors as execution anomalies. Diagnosing the root cause of execution anomalies is critical for ensuring reliable, safe deployment.
Execution anomalies generally arise from two broad sources: (1) agent errors: systemic errors in agent modeling or training, which results in an incorrect policy, or (2) environment errors: unfavorable environment configuration that makes task success inherently infeasible, even for an ideal agent. Agent errors may arise from model defects in the form of inaccuracies in the state representation, reward function, or both, in hand-crafted or learned model used for decision-making (Amodei et al. 2016; Hadfield-Menell et al. 2017; Saisubramanian, Kamar, and Zilberstein 2020); or training flaws in model-free settings, such as suboptimal choices of learning algorithms or sim-to-real gaps (Ramakrishnan et al. 2020). On the other hand, some environment configurations are intrinsically unfavorable to agent success, such as poorly placed air vents in warehouses that reduce the efficiency of robot navigation (Simon 2019), and some tasks are inherently infeasible, such as painting the wall in blue color and red color at the same time.
@@
As both environments and agents grow more complex, identifying the source of execution anomalies becomes increasingly difficult. In practice, such anomalies are often incorrectly and reflexively attributed solely to agent errors. However, if the root cause lies in the environment configuration, no amount of training or verification will resolve the issue unless the environment itself is modified. Without a principled investigation that involves checking for alternative feasible paths using a model-independent planner or simulating task variants, it is difficult to determine whether the issue lies in the agent or the environment. While prior works have focused on testing for model errors (He et al. 2024; Nayyar, Verma, and Srivastava 2022; Pang, Yuan, and Wang 2022) or using formal verification methods to provide guarantees on the occurrence of anomalies (Corsi, Marchesini, and Farinelli 2021; Shea-Blymyer and Abbas 2024), they do not determine whether an anomaly is due to the agent or the environment, without requiring detailed internal access to the agent.
@@
By comparing agent behavior with planner performance across a large suite of environment-task instances, AIProbe can automatically detect model flaws, environment-induced infeasibility, and edge-case behaviors, thereby identifying the operational boundary of safe deployment of autonomous systems. This analysis is also critical for reliability assessment (Olamide, Kuyoro‘Shade, and Oludele 2020) and to generate model cards for autonomous systems (Mitchell et al. 2019). Our evaluation on five domains shows that our black-box differential testing method outperforms the state-of-the-art methods in error detection, in both discrete and continuous settings.
@@
When an agent is unable to reach the goal, we want to automatically distinguish between the cases represented in Figures 4(a) and 4(b). As agent architectures grow more complex, especially with learned models, they end up being considered as a black-box model by the evaluators (Nayyar, Verma, and Srivastava 2022). Error detection in a black-box system deployed in large, complex environments is particularly challenging, as we cannot directly inspect if the agent’s model accounts for lava states or manually analyze the environment-task configuration.
@@
We compare the performance of AIProbe, with 10 and 20 seeds used to generate environment-task configurations, with that of two state-of-the-art fuzz testing approaches designed specifically to test autonomous systems: MDPFuzz (Pang, Yuan, and Wang 2022) and CureFuzz (He et al. 2024). We also perform two ablation studies on components that are critical to AIProbe: environment-task generation and baseline planner. Specifically, we compare AIProbe’s performance to using GPT-4o for environment-task generation, and our proposed heuristic-search baseline planner with that of Breadth First Search (BFS).
@@
It is practically infeasible to manually evaluate an autonomous system on all possible scenarios it may encounter in the deployed environment, which motivated automating the testing process (Karimoddini et al. 2022). Fuzz testing is a software testing technique that involves testing the system on a large number of random inputs to uncover bugs, crashes, security vulnerabilities, or other unexpected behavior. Prior works that apply fuzz testing to autonomous agents, such as CureFuzz (He et al. 2024) and MDPFuzz (Pang, Yuan, and Wang 2022), focus on generating diverse input scenarios, using techniques such as curiosity-driven exploration to uncover edge cases that lead to agent failure. Another line of work uses search-based methods to generate adversarial test cases in the form of environment-task configurations where the agent will likely fail, which can be integrated with fuzz-testing (Tappler et al. 2022, 2024). The search-based methods require access to the agent’s internal model to generate test cases and therefore cannot be applied to black-box systems. Alternatively, differential testing has also been used to evaluate model updates to the agent (Nayyar, Verma, and Srivastava 2022). All these approaches primarily focus on detecting model-specific failures without explicitly addressing the feasibility of tasks within the environment itself. Further, many of them require access to the agent’s internal model. In contrast, AIProbe can detect both environment errors and agent errors in black-box systems.

### v2

@@ Introduction
Autonomous agents are increasingly deployed in complex real-world applications such as autonomous driving (Yurtsever et al. 2020), crop fertilization (Gautron et al. 2022; Solow et al. 2025), and elderly care (Bardaro et al. 2022; Mhlanga 2024). Agents operating in complex settings may sometimes produce undesirable behaviors, including failure to complete the task. We refer to such behaviors as execution anomalies. Diagnosing the root cause of execution anomalies is critical for ensuring reliable, safe deployment.
Execution anomalies generally arise from two broad sources: (1) agent errors: systemic errors in agent modeling or training, which results in an incorrect policy, or (2) environment errors: unfavorable environment configuration that makes task success inherently infeasible, even for an ideal agent. Agent errors may arise from model defects in the form of inaccuracies in the state representation, reward function, or both, in hand-crafted or learned model used for decision-making (Amodei et al. 2016; Hadfield-Menell et al. 2017; Saisubramanian et al. 2020); or training flaws in model-free settings, such as suboptimal choices of learning algorithms or sim-to-real gaps (Ramakrishnan et al. 2020). On the other hand, some environment configurations are intrinsically unfavorable to agent success, such as poorly placed air vents in warehouses that reduce the efficiency of robot navigation (Simon 2019), and some tasks are inherently infeasible, such as painting the wall in blue color and red color at the same time.
@@
As both environments and agents grow more complex, identifying the source of execution anomalies becomes increasingly difficult. In practice, such anomalies are often incorrectly and reflexively attributed solely to agent errors. However, if the root cause lies in the environment configuration, no amount of training or verification will resolve the issue unless the environment itself is modified. Without a principled investigation that involves checking for alternative feasible paths using a model-independent planner or simulating task variants, it is difficult to determine whether the issue lies in the agent or the environment. While prior works have focused on testing for model errors (He et al. 2024; Nayyar et al. 2022; Pang et al. 2022) or using formal verification methods to provide guarantees on the occurrence of anomalies (Corsi et al. 2021; Shea-Blymyer and Abbas 2024), they do not determine whether an anomaly is due to the agent or the environment, without requiring detailed internal access to the agent.
@@
By comparing agent behavior with planner performance across a large suite of environment-task instances, AIProbe can automatically detect model flaws, environment-induced infeasibility, and edge-case behaviors, thereby identifying the operational boundary of safe deployment of autonomous systems. This analysis is also critical for reliability assessment (Olamide et al. 2020) and to generate model cards for autonomous systems (Mitchell et al. 2019). Our evaluation on five domains shows that our black-box differential testing method outperforms the state-of-the-art methods in error detection, in both discrete and continuous settings.
@@
When an agent is unable to reach the goal, we want to automatically distinguish between the cases represented in Figures 4(a) and 4(b). As agent architectures grow more complex, especially with learned models, they end up being considered as a black-box model by the evaluators (Nayyar et al. 2022). Error detection in a black-box system deployed in large, complex environments is particularly challenging, as we cannot directly inspect if the agent’s model accounts for lava states or manually analyze the environment-task configuration.
@@
We compare the performance of AIProbe, with 10 and 20 seeds used to generate environment-task configurations, with that of two state-of-the-art fuzz testing approaches designed specifically to test autonomous systems: MDPFuzz (Pang et al. 2022) and CureFuzz (He et al. 2024). We also perform two ablation studies on components that are critical to AIProbe: environment-task generation and baseline planner. Specifically, we compare AIProbe’s performance to using GPT-4o for environment-task generation, and our proposed heuristic-search baseline planner with that of Breadth First Search (BFS).
@@
It is practically infeasible to manually evaluate an autonomous system on all possible scenarios it may encounter in the deployed environment, which motivated automating the testing process (Karimoddini et al. 2022). Fuzz testing is a software testing technique that involves testing the system on a large number of random inputs to uncover bugs, crashes, security vulnerabilities, or other unexpected behavior. Prior works that apply fuzz testing to autonomous agents, such as CureFuzz (He et al. 2024) and MDPFuzz (Pang et al. 2022), focus on generating diverse input scenarios, using techniques such as curiosity-driven exploration to uncover edge cases that lead to agent failure. Another line of work uses search-based methods to generate adversarial test cases in the form of environment-task configurations where the agent will likely fail, which can be integrated with fuzz-testing (Tappler et al. 2022, 2024). The search-based methods require access to the agent’s internal model to generate test cases and therefore cannot be applied to black-box systems. Alternatively, differential testing has also been used to evaluate model updates to the agent (Nayyar et al. 2022). All these approaches primarily focus on detecting model-specific failures without explicitly addressing the feasibility of tasks within the environment itself. Further, many of them require access to the agent’s internal model. In contrast, AIProbe can detect both environment errors and agent errors in black-box systems.